# Test-Time 3D Occupancy Prediction

### v3

@@ 3.2 Track Dynamic Gaussians
$\operatorname{GaussianOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(Arxiv^{\prime}24)}}}$ [GaussianOcc] C 11.26 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29
@@
$\operatorname{GaussianTR}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}25)}}}$ [GaussianTR] C 11.70 2.09 5.22 14.07 20.43 5.70 7.08 5.12 3.93 0.92 13.36 39.44 15.68 22.89 21.17 21.87
@@
$\operatorname{LangOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(3DV^{\prime}25)}}}$ [LangOcc] C 11.84 3.10 9.00 6.30 14.20 0.40 10.80 6.20 9.00 3.80 10.70 43.70 9.50 26.40 19.60 26.40
@@
$\operatorname{VEON-LiDAR}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ECCV^{\prime}24)}}}$ [VEON] C&L 15.14 10.40 6.20 17.70 12.70 8.50 7.60 6.50 5.50 8.20 11.80 54.50 25.50 30.20 25.40 25.40
@@
$\operatorname{SelfOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}24)}}}$ [SelfOcc] C 10.54 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60
@@ 3.3 Gaussian Voxelization
Following 3DGS [3DGS], our model refines Gaussian parameters at test time by minimizing a loss that enforces color consistency, with sky regions intentionally masked out as in [DrivingGaussian]. Optionally, to further mitigate the inherent noise and errors in VFMs’ predictions and tracking results, we introduce a Trilateral Radial Basis Function (TRBF) kernel for periodic smoothing and denoising. TRBF kernel improves the spatial and temporal coherence of occupancy predictions by leveraging spatial, radiometric, and semantic affinities among Gaussians for anisotropic information propagation while preserving local structures and semantic boundaries. Formally, for each $\mathbf{m}_{i}\in\mathbf{G}_{i}^{(t)}$, the kernel smoothing is defined as a deformable convolution over its nearest neighbors:
@@
where each term $\operatorname{attr}\in\{\bm{\mu},\mathbf{c},\mathbf{m}\}$ is defined as the following format,
@@
For efficient occupancy estimation, we voxelize the accumulated Gaussians $\mathcal{G}^{(t)}$ into a discrete grid $\mathbf{\Omega}=[\frac{X}{\delta}\times\frac{Y}{\delta}\times\frac{Z}{\delta}]$, where each Gaussian’s contribution on a voxel is weighted based on its spatial proximity. Formally, the semantic probability of a voxel $v\in\mathbf{\Omega}$ is given by,
@@
Table 2: 3D occupancy prediction performance on the high-resolution nuCraft dataset [nuCraft]. Pretraining time is reported in GPU hours.
@@
$\operatorname{C-CONet}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ICCV^{\prime}23)}}}$ [C-CONet] Dense - 13.4 14.30 9.10 16.50 18.30 7.40 12.30 11.10 9.40 5.80 13.20 32.50 - - - 19.90
$\operatorname{SelfOcc}^{\dagger}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}24)}}}$ [SelfOcc] 384 hrs 2.22 0.41 0.54 2.79 7.12 0.00 0.81 1.67 0.00 0.00 5.50 2.41 3.88 3.55 1.96 2.72
@@ 4.1 Experiment Setup
Experiments were conducted on the widely used nuScenes [nuScenes] benchmark using 3D occupancy GT from Occ3D-nuScenes [Occ3D] and nuCraft [nuCraft]. The nuScenes dataset consists of 600 training scenes and 150 validation ones. Existing supervised and self-supervised methods typically require extensive offline training on the training split. In contrast, TT-Occ requires no pretraining and is directly evaluated on the validation split. In particular, Occ3D-nuScenes [Occ3D] provides voxelized occupancy annotations at 0.4m resolution, covering a spatial range o [40m,40m][-40m,40m] along the X and Y axes and [1m,5.4m][-1m,5.4m] along the Z axis. nuCraft [nuCraft] offers more finer-grained annotations with a resolution of 0.2m, covering [51.2m,51.2m][-51.2m,51.2m] in the X and Y directions and [5m,3m][-5m,3m] in the Z direction.
We evaluate semantic occupancy prediction using mean Intersection over Union ($\operatorname{mIoU}$), computed as the average $\operatorname{IoU}$ across all classes. Following prior works [OccNeRF, GaussianOcc, GaussianTR], we exclude the “noise” and “other flat” categories, as these do not correspond to valid prompts in open-vocabulary segmentation. We primarily compare our method with self-supervised counterparts, including SimpleOcc [SimpleOcc], OccNeRF [OccNeRF], DistillNeRF [DistillNeRF], GaussianOcc [GaussianOcc], GaussianTR [GaussianTR], LangOcc [LangOcc], VEON [VEON], and SelfOcc [SelfOcc]. These methods represent a broad range of self-supervised occupancy research and include both NeRF [NeRF] and 3DGS [3DGS] representation. For reference, we also include results from self-supervised methods that serve as upper bounds for performance comparison. These include BEVFormer [BEVFormer] and CTF-Occ [Occ3D], which are trained with dense 3D voxel-level annotations, as well as RenderOcc [RenderOcc] and OccFlowNet [OccFlowNet], which are trained on sparse point-level ground truth.
Figure 4: Qualitative comparisons on nuCraft [nuCraft] between both variants of the proposed TT-Occ and SelfOcc [SelfOcc].
Table 3: Comparison of TT-Occ and SelfOcc [SelfOcc] using the same OpenSeed [OpenSeeD] VFM under different RayIoU thresholds.
@@ 4.2 Main Results
Results on Occ3D-nuScenes are shown in Table 1. It is evident that both variants of TT-Occ not only eliminate the need for costly offline training but also surpass the previous SOTA. Notably, $\operatorname{TT\text{-}OccLiDAR}$ even achieves an mIoU of 27.41, surpassing RenderOcc [RenderOcc] (23.93), which is trained with sparse 3D ground truth. Moreover, our camera-only variant $\operatorname{TT\text{-}OccCamera}$ reaches an mIoU of 16.70, exceeding VEON-LiDAR [VEON] (15.14), which is trained using LiDAR supervision. In addition, compared to the SOTA self-supervised baseline SelfOcc [SelfOcc], our approach achieves higher IoU not only for frequently occurring, large-area categories such as terrain and vegetation, but also shows substantial improvements on rare, dynamic, and small-area categories such as motorcycle, bus, and pedestrian.
To further evaluate the geometric quality beyond conventional mIoU, we report RayIoU under multiple thresholds, following the standard protocol in [rayiou]. Importantly, all methods are equipped with the same OpenSeed [OpenSeeD] VFM for fair comparison. As shown in Table 3, both TT-Occ variants achieve consistent gains across all RayIoU levels, indicating that our time-aware 3D Gaussians improve not only occupancy classification but also the metric accuracy of reconstructed geometry along camera rays. TT-OccCamera improves the RayIoU@4 metric by 30.8% over SelfOcc, while TT-OccLiDAR delivers the strongest performance with a substantial 115% improvement. These results highlight that TT-Occ suppresses temporal drift and produces cleaner free-space estimates, where those qualities that are not fully captured by mIoU alone. In practice, higher RayIoU translates to more reliable scene geometry for downstream planning and safety-critical perception modules, underscoring the benefit of test-time Gaussian modeling even when using identical VFM semantics.
@@
Results on nuCraft are summarized in Table 2. As no prior self-supervised methods have been trained or evaluated under this setting, we adapt SelfOcc [SelfOcc] using its official implementation and checkpoint as a baseline for comparison. As shown in the table, TT-Occ consistently and significantly outperforms SelfOcc, demonstrating superior adaptability and robustness across varying resolutions.
Qualitative comparisons on nuCraft between both variants of TT-Occ and SelfOcc [SelfOcc] are shown in Fig. 4. Several key observations emerge from these results. (1) Both our LiDAR- and camera-based variants produce highly accurate occupancy predictions that closely align with the ground truth. In contrast, SelfOcc generates overly dense predictions, assigning occupancy to nearly all voxels, including empty regions. This not only incurs significant computational redundancy but also results in severe discrepancies with the ground truth, particularly around dynamic objects (see the radial blue regions). (2) The LiDAR-based variant produces geometrically accurate reconstructions with broad spatial coverage. However, its fidelity is inherently constrained by the sparsity of LiDAR returns, especially for small or partially scanned objects such as vehicles. (3) The camera-based variant offers denser reconstructions and better captures small objects within the field of view. Nonetheless, it may struggle with distant regions due to occlusions or limited depth resolution, and the geometry inferred from depth estimation is generally less accurate than that derived from LiDAR. Despite these challenges, $\operatorname{TT\text{-}OccCamera}$ still remains the state-of-the-art among vision-only occupancy methods. Moreover, thanks to the modular design of our system, it can be readily enhanced by integrating more advanced VFMs, and thus continues to benefit from the rapid progress in this area.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of each component in TT-Occ, we conduct ablation studies on a 10% held-out validation subset of Occ3D-nuScenes [Occ3D]. Since dynamic classes typically occupy only a small portion of the scene but play a critical role in both human perception and downstream tasks, we report not only the overall IoU and mIoU, but also the IoU of representative dynamic classes (bus, pedestrian) and a representative static class (manmade). We use 3DGS [3DGS] as the baseline ($\mathbb{A}$), where Gaussians are initialized using the “lift” strategy introduced in Section 3.1 at each time step without temporal information. Gaussians are voxelized by directly scattering their centers. As shown in the tables present in Fig. 5, this naïve approach yields poor results due to sparse observations, emphasizing the importance of using anisotropic Gaussian occupancy to better approximate scene geometry. Next, we introduce covariance-aware voxelization (Eq. (6)) and apply sigmoid-based scale regulation ($\mathbb{B}$). These lead to consistent improvements across both static and dynamic classes for both LiDAR and camera inputs.
@@
Both Variants $\mathbb{A}$ and $\mathbb{B}$ are single-frame models. Allowing Gaussians to accumulate across frames ($\mathbb{C}$) greatly improves the overall and static class performance (e.g., manmade) due to the aggregation of Gaussians for static content, which dominates the scene. However, dynamic class performance drops significantly, as untracked accumulation of moving Gaussians causes temporal inconsistency (see $\mathbb{C}$ in Fig. 3 for trailing artifacts). To address this, we incorporate tracking dynamic Gaussians as described in Section 3.2, which significantly improves the accuracy of dynamic classes while maintaining performance on static content. As shown in $\mathbb{D}$, this yields cleaner occupancy with trailing and ghosting artifacts largely eliminated. The ablation study on the optional TRBF fusion module is presented in Appendix.
Table 4: Impact of semantic and geometric VFMs of $\mathrm{TT\text{-}Occ}$ on the full validation set of Occ3D-nuScenes.
Method Semantic Model Geometry mIoU $\uparrow$
$\mathrm{SelfOcc}$ $\mathrm{OpenSeeD}$ [OpenSeeD] 11.6
$\mathrm{TT\text{-}OccLiDAR}$ $\mathrm{OpenSeeD}$ [OpenSeeD] 23.6
$\mathrm{GroundedSAM2}$ [groundedsam] 21.3
$\mathrm{REX\text{-}Omni}$ [rexomni] 27.4
$\mathrm{TT\text{-}OccCamera}$ $\mathrm{OpenSeeD}$ [OpenSeeD] $\mathrm{VGGT}$ [vggt] 13.4
$\mathrm{OpenSeeD}$ [OpenSeeD] $\mathrm{MapAnything}$ [keetha2025mapanything] 14.3
$\mathrm{GroundedSAM2}$ [groundedsam] $\mathrm{VGGT}$ [vggt] 12.2
$\mathrm{REX\text{-}Omni}$ [rexomni] $\mathrm{VGGT}$ [vggt] 15.9
$\mathrm{REX\text{-}Omni}$ [rexomni] $\mathrm{MapAnything}$ [keetha2025mapanything] 16.7
@@
Ablation on Loosely Coupled VFM Components. To demonstrate that our framework is loosely coupled and can support arbitrary VFMs, and to further investigate how different VFMs affect overall performance, we integrate three semantic models, namely OpenSeeD [OpenSeeD], GroundingSAM2 [groundedsam], and REX-Omni [rexomni], as well as two depth-estimation backbones, VGGT [vggt] and MapAnything [keetha2025mapanything]. The results of all combinations are summarized in Table 4. It can be seen that GroundingSAM2 yields the weakest performance, while REX-Omni provides the strongest semantic cues. Among the depth models, MapAnything consistently outperforms VGGT, primarily because it provides metric-scale depth directly, whereas VGGT requires additional triangulation to recover metric scale, which inevitably introduces errors. Overall, these experiments confirm that our modular system design can directly benefit from the rapid progress of VFMs.
@@
Efficiency Analysis. We provide a detailed runtime breakdown of our pipeline for the vision-centric and LiDAR-based variants in Fig. 6 (using OpenSeeD [OpenSeeD]). The reported values represent the average processing time per timestep across six input images. Semantic segmentation using OpenSeeD [OpenSeeD] constitutes the most computationally intensive step in both pipelines, accounting for 28.5% of total runtime in the camera variant and 77.9% in the LiDAR variant. In the vision-centric scenario, the absence of LiDAR data requires additional processes such as depth estimation, triangulation-based calibration, and point cloud denoising, collectively contributing 46.5% of the overall runtime. Gaussian voxelization and the optional TRBF fusion module are relatively efficient; however, their runtime is proportional to the number of Gaussians involved. Therefore, the camera-based pipeline which has denser Gaussians experiences slightly increased computational overhead compared to its LiDAR-based counterpart. Finally, tracking dynamic Gaussians in $\operatorname{TT\text{-}OccCamera}$ incurs much higher computational costs compared to $\operatorname{TT\text{-}OccLiDAR}$ due to its reliance on dense optical flow estimation across six images using RAFT [raft], whereas the LiDAR variant only applies lightweight ICP alignment for sparse foreground points.
@@
Figure 6: Inference time comparison among $T$-OccCamera, $T$-OccLiDAR, and SelfOcc [SelfOcc]. The horizontal stacked bars illustrate the per-module runtime composition of each module.
@@
We report the average peak GPU memory usage on high-resolution nuCraft [nuCraft] in Table 5. Because our system processes modules sequentially, the peak memory footprint is determined by the most demanding component at runtime (e.g., OpenSeeD or the accumulated Gaussians). In the LiDAR variant, memory usage is primarily dominated by the semantic module, whereas in the camera variant, the growing set of Gaussians becomes the major contributor. Overall, TT-Occ maintains a peak memory consumption below 10GB for both LiDAR and camera settings.
@@
$T$-OccLiDAR 786,874 8.5 (dom. by OpenSeeD)
$T$-OccCamera 2,968,647 9.9 (dom. by Gaussians)

### v4

@@ 3.2 Track Dynamic Gaussians
$\operatorname{GaussianOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(Arxiv^{\prime}24)}}}$ [10] C 11.26 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29
@@
$\operatorname{GaussianTR}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}25)}}}$ [16] C 11.70 2.09 5.22 14.07 20.43 5.70 7.08 5.12 3.93 0.92 13.36 39.44 15.68 22.89 21.17 21.87
@@
$\operatorname{LangOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(3DV^{\prime}25)}}}$ [3] C 11.84 3.10 9.00 6.30 14.20 0.40 10.80 6.20 9.00 3.80 10.70 43.70 9.50 26.40 19.60 26.40
@@
$\operatorname{VEON-LiDAR}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ECCV^{\prime}24)}}}$ [53] C&L 15.14 10.40 6.20 17.70 12.70 8.50 7.60 6.50 5.50 8.20 11.80 54.50 25.50 30.20 25.40 25.40
@@
$\operatorname{SelfOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}24)}}}$ [12] C 10.54 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60
@@ 3.3 Gaussian Voxelization
Following 3DGS [20], our model refines Gaussian parameters at test time by minimizing a loss that enforces color consistency, with sky regions intentionally masked out as in [56]. Optionally, to further mitigate the inherent noise and errors in VFMs’ predictions and tracking results, we introduce a Trilateral Radial Basis Function (TRBF) kernel for periodic smoothing and denoising. TRBF kernel improves the spatial and temporal coherence of occupancy predictions by leveraging spatial, radiometric, and semantic affinities among Gaussians for anisotropic information propagation while preserving local structures and semantic boundaries. Formally, for each $\mathbf{m}_{i}\in\mathbf{G}_{i}^{(t)}$, the kernel smoothing is defined as a deformable convolution over its nearest neighbors:
@@
where each term $\operatorname{attr}\in\{\bm{\mu},\mathbf{c},\mathbf{m}\}$ is defined as the following format:
@@
For efficient occupancy estimation, we voxelize the accumulated Gaussians $\mathcal{G}^{(t)}$ into a discrete grid $\mathbf{\Omega}=[\frac{X}{\delta}\times\frac{Y}{\delta}\times\frac{Z}{\delta}]$, where each Gaussian’s contribution on a voxel is weighted based on its spatial proximity. Formally, the semantic probability of a voxel $v\in\mathbf{\Omega}$ is given by:
@@
Table 2: 3D occupancy prediction performance on the high-resolution nuCraft dataset [58]. Pretraining time is reported in GPU hours.
@@
$\operatorname{C-CONet}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ICCV^{\prime}23)}}}$ [41] Dense - 13.4 14.30 9.10 16.50 18.30 7.40 12.30 11.10 9.40 5.80 13.20 32.50 - - - 19.90
$\operatorname{SelfOcc}^{\dagger}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(CVPR^{\prime}24)}}}$ [12] 384 hrs 2.22 0.41 0.54 2.79 7.12 0.00 0.81 1.67 0.00 0.00 5.50 2.41 3.88 3.55 1.96 2.72
@@ 4.1 Experiment Setup
Experiments were conducted on the widely used nuScenes [5] benchmark using 3D occupancy GT from Occ3D-nuScenes [35] and nuCraft [58]. The nuScenes dataset consists of 600 training scenes and 150 validation ones. Existing supervised and self-supervised methods typically require extensive offline training on the training split. In contrast, TT-Occ requires no pretraining and is directly evaluated on the validation split. In particular, Occ3D-nuScenes [35] provides voxelized occupancy annotations at 0.4m resolution, covering a spatial range o [40m,40m][-40m,40m] along the X and Y axes and [1m,5.4m][-1m,5.4m] along the Z axis. nuCraft [58] offers more finer-grained annotations with a resolution of 0.2m, covering [51.2m,51.2m][-51.2m,51.2m] in the X and Y directions and [5m,3m][-5m,3m] in the Z direction.
We evaluate semantic occupancy prediction using mean Intersection over Union ($\operatorname{mIoU}$), computed as the average $\operatorname{IoU}$ across all classes. Following prior works [50, 10, 16], we exclude the “noise” and “other flat” categories, as these do not correspond to valid prompts in open-vocabulary segmentation. We primarily compare our method with self-supervised counterparts, including SimpleOcc [11], OccNeRF [50], DistillNeRF [39], GaussianOcc [10], GaussianTR [16], LangOcc [3], VEON [53], and SelfOcc [12]. These methods represent a broad range of self-supervised occupancy research and include both NeRF [26] and 3DGS [20] representation. For reference, we also include results from self-supervised methods that serve as upper bounds for performance comparison. These include BEVFormer [23] and CTF-Occ [35], which are trained with dense 3D voxel-level annotations, as well as RenderOcc [27] and OccFlowNet [4], which are trained on sparse point-level ground truth. All experiments were conducted on an NVIDIA RTX 6000 Ada (48GB) GPU. Please refer to the supplementary materials for implementation details.
Figure 4: Qualitative comparisons on nuCraft [58] between both variants of the proposed TT-Occ and SelfOcc [12].
Table 3: Comparison of TT-Occ and SelfOcc [12] using the same OpenSeed [51] VFM under different RayIoU thresholds.
@@ 4.2 Main Results
Results on Occ3D-nuScenes are shown in Table 1. It is evident that both variants of TT-Occ not only eliminate the need for costly offline training but also surpass the previous SOTA. Notably, $\operatorname{TT\text{-}OccLiDAR}$ even achieves an mIoU of 27.41, surpassing RenderOcc [27] (23.93), which is trained with sparse 3D ground truth. Moreover, our camera-only variant $\operatorname{TT\text{-}OccCamera}$ reaches an mIoU of 16.70, exceeding VEON-LiDAR [53] (15.14), which is trained using LiDAR supervision. In addition, compared to the SOTA self-supervised baseline SelfOcc [12], our approach achieves higher IoU not only for frequently occurring, large-area categories such as terrain and vegetation, but also shows substantial improvements on rare, dynamic, and small-area categories such as motorcycle, bus, and pedestrian.
To further evaluate the geometric quality beyond conventional mIoU, we report RayIoU under multiple thresholds, following the standard protocol in [24]. Importantly, all methods are equipped with the same OpenSeed [51] VFM for fair comparison. As shown in Table 3, both TT-Occ variants achieve consistent gains across all RayIoU levels, indicating that our time-aware 3D Gaussians improve not only occupancy classification but also the metric accuracy of reconstructed geometry along camera rays. TT-OccCamera improves the RayIoU@4 metric by 30.8% over SelfOcc, while TT-OccLiDAR delivers the strongest performance with a substantial 115% improvement. These results highlight that TT-Occ suppresses temporal drift and produces cleaner free-space estimates, where those qualities that are not fully captured by mIoU alone. In practice, higher RayIoU translates to more reliable scene geometry for downstream planning and safety-critical perception modules, underscoring the superiority of TT-Occ given the same semantic VFM.
@@
Results on nuCraft are summarized in Table 2. As no prior self-supervised methods have been trained or evaluated under this setting, we adapt SelfOcc [12] using its official implementation and checkpoint as a baseline for comparison. As shown in the table, TT-Occ consistently and significantly outperforms SelfOcc, demonstrating superior adaptability and robustness across varying resolutions.
Qualitative comparisons on nuCraft between both variants of TT-Occ and SelfOcc [12] are shown in Fig. 4. Several key observations emerge from these results. (1) Both our LiDAR and camera variants produce highly accurate occupancy predictions that closely align with the ground truth. In contrast, SelfOcc generates overly dense predictions, assigning occupancy to nearly all voxels, including empty regions. This not only incurs significant computational redundancy but also results in severe discrepancies with the ground truth, particularly around dynamic objects (see the radial blue regions). (2) The LiDAR-based variant produces geometrically accurate reconstructions with broad spatial coverage. However, its fidelity is inherently constrained by the sparsity of LiDAR returns, especially for small or partially scanned objects such as vehicles. (3) The camera-based variant offers denser reconstructions and better captures small objects within the field of view. Nonetheless, it may struggle with distant regions due to occlusions or limited depth resolution, and the geometry inferred from depth estimation is generally less accurate than that derived from LiDAR. Despite these challenges, $\operatorname{TT\text{-}OccCamera}$ still remains the state-of-the-art among vision-only occupancy methods.
@@ 4.3 Ablation Studies
To evaluate the effectiveness of each component in TT-Occ, we conduct ablation studies on a 10% held-out validation subset of Occ3D-nuScenes [35]. Since dynamic classes typically occupy only a small portion of the scene but play a critical role in both human perception and downstream tasks, we report not only the overall IoU and mIoU, but also the IoU of representative dynamic classes (bus, pedestrian) and a representative static class (manmade). We use 3DGS [20] as the baseline ($\mathbb{A}$), where Gaussians are initialized using the “lift” strategy introduced in Section 3.1 at each time step without temporal information. Gaussians are voxelized by directly scattering their centers. As shown in the tables present in Fig. 5, this naïve approach yields poor results due to sparse observations, emphasizing the importance of using anisotropic Gaussian occupancy to better approximate scene geometry. Next, we introduce covariance-aware voxelization (Eq. (6)) and apply sigmoid-based scale regulation ($\mathbb{B}$). These lead to consistent improvements across both static and dynamic classes for both LiDAR and camera inputs.
@@
Both Variants $\mathbb{A}$ and $\mathbb{B}$ are single-frame models. Allowing Gaussians to directly accumulate across frames without tracking ($\mathbb{C}$) greatly improves the overall and static class performance (e.g., manmade) due to the aggregation of Gaussians for static content, which dominates the scene. However, dynamic class performance drops significantly, as untracked accumulation of moving Gaussians causes temporal inconsistency (see $\mathbb{C}$ in Fig. 3 for trailing artifacts). To address this, we incorporate tracking dynamic Gaussians as described in Section 3.2, which significantly improves the accuracy of dynamic classes while maintaining performance on static content. As shown in $\mathbb{D}$, this yields cleaner occupancy with trailing and ghosting artifacts largely eliminated. The ablation study on the optional TRBF fusion module is presented in supplementary materials.
Table 4: Impact of semantic and geometric VFMs of $\mathrm{TT\text{-}Occ}$.
Method Semantic Geometry mIoU $\uparrow$
$\mathrm{SelfOcc}$ $\mathrm{OpenSeeD}$ [51] 11.6
$\mathrm{TT\text{-}OccLiDAR}$ $\mathrm{OpenSeeD}$ [51] 23.6
$\mathrm{GroundedSAM2}$ [30] 21.3
$\mathrm{REX\text{-}Omni}$ [17] 27.4
$\mathrm{TT\text{-}OccCamera}$ $\mathrm{OpenSeeD}$ [51] $\mathrm{VGGT}$ [37] 13.4
$\mathrm{OpenSeeD}$ [51] $\mathrm{MapAnything}$ [19] 14.3
$\mathrm{GroundedSAM2}$ [30] $\mathrm{VGGT}$ [37] 12.2
$\mathrm{REX\text{-}Omni}$ [17] $\mathrm{VGGT}$ [37] 15.9
$\mathrm{REX\text{-}Omni}$ [17] $\mathrm{MapAnything}$ [19] 16.7
@@
Ablation on Loosely Coupled VFM Components. To demonstrate that our framework is loosely coupled and can support arbitrary VFMs, and to further investigate how different VFMs affect overall performance, we integrate three semantic models, namely OpenSeeD [51], GroundingSAM2 [30], and REX-Omni [17], as well as two depth-estimation backbones, VGGT [37] and MapAnything [19]. Results of all combinations on Occ3D-nuScenes [35] are summarized in Table 4. GroundingSAM2 yields the weakest performance, while REX-Omni provides the strongest semantic cues. Among depth models, MapAnything consistently outperforms VGGT, primarily because it provides metric-scale depth directly, whereas VGGT requires additional triangulation to recover metric scale, which inevitably introduces errors. Overall, these experiments confirm that our loosely coupled, modular design allows seamless integration of advanced VFMs without modification, directly benefiting from their rapid evolution and scaling trend.
@@
Efficiency Analysis. We provide a detailed runtime breakdown of our pipeline for the vision-centric and LiDAR-based variants in Fig. 6 (using OpenSeeD [51]). The reported values represent the average processing time per timestep across six input images. Semantic segmentation using OpenSeeD [51] constitutes the most computationally intensive step in both pipelines, accounting for 28.5% of total runtime in the camera variant and 77.9% in the LiDAR variant. In the vision-centric scenario, the absence of LiDAR data requires additional processes such as depth estimation, triangulation-based calibration, and point cloud denoising, collectively contributing 46.5% of the overall runtime. Gaussian voxelization and the optional TRBF fusion module are relatively efficient; however, their runtime is proportional to the number of Gaussians involved. Therefore, the camera-based pipeline which has denser Gaussians experiences slightly increased computational overhead compared to its LiDAR-based counterpart. Finally, tracking dynamic Gaussians in $\operatorname{TT\text{-}OccCamera}$ incurs much higher computational costs compared to $\operatorname{TT\text{-}OccLiDAR}$ due to its reliance on dense optical flow estimation across six images using RAFT [34], whereas the LiDAR variant only applies lightweight ICP alignment for sparse foreground points.
@@
Figure 6: Inference time comparison among $T$-OccCamera, $T$-OccLiDAR, and SelfOcc [12]. The horizontal stacked bars illustrate the per-module runtime composition of each module.
@@
We report the average peak GPU memory usage on high-resolution nuCraft [58] in Table 5. Because our system processes modules sequentially, the peak memory footprint is determined by the most demanding component at runtime (e.g., OpenSeeD or the accumulated Gaussians). In the LiDAR variant, memory usage is primarily dominated by the semantic module, whereas in the camera variant, the growing set of Gaussians becomes the major contributor. Overall, TT-Occ maintains a peak memory consumption below 10GB for both LiDAR and camera settings.
@@
$\mathrm{TT\text{-}OccLiDAR}$ 786,874 5.6 (dom. by OpenSeeD)
$\mathrm{TT\text{-}OccCamera}$ 2,968,647 9.9 (dom. by Gaussians)